# State Space Models for Event Cameras

### v2

@@ Abstract
Today, state-of-the-art deep neural networks that process event-camera data first convert a temporal window of events into dense, grid-like input representations. As such, they exhibit poor generalizability when deployed at higher inference frequencies (i.e., smaller temporal windows) than the ones they were trained on. We address this challenge by introducing state-space models (SSMs) with learnable timescale parameters to event-based vision. This design adapts to varying frequencies without the need to retrain the network at different frequencies. Additionally, we investigate two strategies to counteract aliasing effects when deploying the model at higher frequencies. We comprehensively evaluate our approach against existing methods based on RNN and Transformer architectures across various benchmarks, including Gen1 and 1 Mpx event camera datasets. Our results demonstrate that SSM-based models train 33% faster and also exhibit minimal performance degradation when tested at higher frequencies than the training input. Traditional RNN and Transformer models exhibit performance drops of more than 20 mAP, with SSMs having a drop of 3.31 mAP, highlighting the effectiveness of SSMs in event-based vision tasks.
Figure 1: Top-Left Previous works [13, 34] use RNN architectures with convolutional or attention mechanisms to train models that have superior performance on downstream tasks. However, the use of RNNs leads to slower training and the learned weights only generalize well to data deployed at the same frequency as that used at training time. Bottom-Left We solve this problem by utilizing SSMs for temporal aggregation, which enables faster training by either utilizing the S4 model [19] or S5 [41] parallel scans. By their nature, these models allow deployment at different frequencies than those used at training time since they have a learnable timescale parameter. Right Our SSM-based models achieve an average performance drop between training and testing frequencies of 3.31 mAP averaged on both Gen1 [8] and 1 Mpx [34] datasets, while RVT [13] and GET [33] have a drop of 21.25 and 24.53 mAP, respectively.
@@ 1 Introduction
State-space models [16] function as CNN during training and are converted to an efficient RNN at test time. To achieve this, S4 [19] employs a technique where the SSM, which is not practical for training on modern hardware due to its sequential nature, is transformed into a CNN by unrolling. S5 [41] uses parallel scans during training, and is employed as RNN during inference. State-space models [16] can be deployed at different frequencies at inference time because they are Linear Time-Invariant (LTI) continuous-time systems that can be transformed into a discrete-time system with an arbitrary step size. This feature permits inference at arbitrary frequencies by globally adjusting the timescale parameter based on the ratio between the new and old frequency sampling rates. Consequently, we tackle the longstanding issue in event-based vision, which requires multiple training cycles with different frequencies to adapt the neural network for various frequencies during inference.
For the task of object detection, we find that incorporating state-space layers accelerates the training by up to 33% relative to existing recurrent vision transformer approaches [13, 29]. This is achieved while maintaining performance comparable to existing methods. Notably, our approach demonstrates superior generalization to higher temporal frequencies with a drop of only 3.31 mAP, while previous methods experience a drop of 21.25 mAP or more. Also, we achieve comparable performance to RVT although we use a linear state-space model rather than a non-linear LSTM model. This also shows that the complexity of the LSTM might not be needed. For this to work, we introduce two strategies (frequency-selective masking and $H_{2}$ norm) to counteract the aliasing effect encountered with increased temporal frequencies in event cameras. First one is a low-pass bandlimiting modification to SSM that encourages the learned convolutional kernels to be smooth. Second one mitigates the aliasing problem by attenuating the frequency response after a chosen frequency. We argue that state-space modeling offers a significant new direction for research in event-based vision, offering promising solutions to the challenges inherent in processing event-based data effectively and efficiently.
@@
  • Our experimental results outperform existing methods at higher frequencies by 20 mAP on average and show 33% increase in the training speed.
@@ 2.1 Object detection with Event Cameras
The second research branch investigates the use of Graph Neural Networks (GNNs) or Spiking Neural Networks (SNNs). GNNs dynamically construct a spatio-temporal graph where new nodes and edges are instantiated by selectively sub-sampling events and identifying pre-existing nodes that are proximate in the space-time continuum [12, 43, 23]. A pivotal challenge lies in architecting the network such that information can disseminate effectively across extensive regions of the space-time volume. This becomes particularly important when dealing with large objects that exhibit slow relative motion to the camera. Moreover, while aggressive sub-sampling is often necessary to achieve low-latency inference, it introduces the risk of omitting important information from the event stream. On the other hand, SNNs [39, 7, 2] transmit information sparsely within the system. Unlike RNNs, SNNs emit spikes only when a threshold is met, making them hard to optimize due to the non-differentiable spike mechanism. Some solutions [32] bypass the threshold, but this sacrifices sparsity in deeper layers. Overall, SNNs remain a challenging area needing more foundational research for optimized performance.
@@ 2.2 Continuous-time Models
In our study, we empirically show that S4, S4D and S5 models achieve results on-par with state-of-the-art when combined with attention mechanisms on complex data. We also identify and address aliasing issues in these models, proposing two corrective strategies. Our work extends the range and robustness of continuous-time models for complex sequence modeling tasks such as object detection for event cameras.
@@ 3 Method
In this section, we firstly formalize the operating mechanism of event cameras and provide a notation used for describing state-space models in the preliminaries (Sec. 3.1). Secondly, we describe our approach of incorporating variants of state-space models as layers within our block (Sec. 3.2). This innovative design solves the problems associated with slow training and the variable frequency inference for event cameras.
@@ 3.1 Preliminaries
State-Space Models (SSMs). Linear State-Space Models (SSMs) form the crucial part of the backbone in our architecture, where we compare S4 [19], S4D [18] and S5 [41] layer variants. Given an input vector $\mathbf{u}(t)\in\mathbb{R}^{U}$, a latent state vector $\mathbf{x}(t)\in\mathbb{R}^{P}$, and an output vector $\mathbf{y}(t)\in\mathbb{R}^{M}$, the governing equations of a continuous-time linear SSM can be mathematically represented as:
@@ 3.2 SSM-ViT block
Regarding the block structure, initially, the input undergoes convolution with a defined kernel of size $k\times k$ with a stride $s$. This operation effectively captures the essential spatial features of the input. Following the convolution operation, the structure introduces a ’Block-SA’ module. This module is pivotal in implementing self-attention mechanisms, but it does so specifically within local windows. The localized nature of the attention in this block ensures a focus on immediate spatial relations, allowing for a detailed representation of close-proximity features.
@@
The final and crucial component of the block structure is the State-Space Model (SSM) block, that is designed to compute both the output and the state in parallel, either by using S4 [19], S4D [18] or S5 [41] model variant. This ensures temporal consistency and a seamless transition of information between consecutive time steps. Efficient computation is crucial since it allows for the faster training than RNN, and timescale parameter on temporal aggregation is important since we can rescale it during inference and deploy at any frequency we want.
@@ 3.3.1 Output Masking
To modulate the frequency spectrum of the model, we define a hyperparameter $\alpha$, which is used for masking of the computed effective frequency $f_{n}$ for each state:
@@
where $\Delta t$ denotes the discrete time-step, and $r$ is the rate at which we train the model, by default it is 1, rate gets halved when deploying at twice the trained frequency etc.
@@
where $C_{n}$ represents the coefficients in the state representation. The selection of $\alpha$ is critical, with $\alpha=1.0$ representing the Nyquist limit under idealized conditions. However, practical considerations and model constraints typically necessitate a lower empirical threshold. Our findings suggest that setting $\alpha=0.5$ yields optimal outcomes for systems with diagonal state matrices, such as in S4D and S5 configurations.
@@ 4 Experiments
We perform ablation studies and systematic evaluations of our proposed models utilizing both the Gen1 [8] and 1 Mpx [34] event-based camera datasets. We assess the model’s proficiency in adapting to unseen inference frequencies during training time on both datasets. We employed two variants for training across both datasets: the base model ViT-SSM-B, alongside its scaled-down derivative ViT-SSM-S (small). Additionally, to study the robustness and generalization capabilities of our architecture, we subject it to empirical testing on the DSEC dataset [14], which was not part of the original training corpus and the visual results are provided in the appendix.
@@ 4.1 Setup
Implementation details. Our models are trained using 32-bit precision arithmetic across 400k iterations, making use of the ADAM optimizer [25] and a OneCycle learning rate schedule [42] that decays linearly from its highest value. We adopt a mixed batching technique that applies standard Backpropagation Through Time (BPTT) to half of the batch samples and Truncated BPTT (TBPTT) to the other half. This technique was introduced in the RVT [13] and ablated in their supplementary material. We found it to be equally effective also for SSMs. For the S4(D) method, details can be found in [19, 18]. As for S5 [41], we utilize an efficient implementation from scratch in PyTorch; original public version of S5 is available in JAX. Data augmentation is carried out through random horizontal flips and zooming operations, both inward and outward. Event representations are formed based on 50 ms time windows that correspond to 20 Hz sampling frequency, divided into T = 10 discrete bins. We incorporate a YOLOX detection head [11] that integrates IOU, class, and regression losses, averaged across both batch and sequence lengths during each optimization phase. Bandlimiting is implemented during training and inference of the SSM model by masking the output matrix C as explained in 3.3.1. In case of $H_{2}$ norm, we add it as term to the loss function of the model. We conduct training on the GEN1 dataset using A100 GPU, employing a batch size of 8 and a sequence length of 21. We use a global learning rate of 2e-4, which is propagated to the SSM components. For the 1 Mpx dataset, we train with a batch size of 12 and a sequence length of 10, using a learning rate of 3.5e-4 across two A100 GPUs.
@@
The 1 Mpx dataset [34] similarly focuses on driving environments but offers recordings at a higher resolution of 720 $\times$ 1280 pixels, captured over multiple months during both day-time and night-time. It contains roughly 15 hours of event data and accumulates around 25 million bounding box labels, spread across three classes: cars, pedestrians, and two-wheelers. Following original evaluation criteria, we eliminate bounding boxes with a side length under 20 pixels and a diagonal less than 60 pixels, while also reducing the input resolution to 640 $\times$ 360 pixels.
@@ 4.2 Benchmark comparisons
S5-ViT-B (ours) Transformer + SSM YOLOX 47.4 8.16 47.2 9.57 18.2
Table 1: Comparisons on test sets of Gen1 and 1 Mpx datasets (20 Hz). Best results in bold and second best underlined. A star $*{}^{*}$ suggests that this information was not directly available and estimated based on the publications. Runtime is measured in milliseconds for a batch size of 1. We used a T4 GPU for SSM-ViT to compare against indicated timings in prior work [34] on comparable GPUs (Titan Xp). Ours achieves comparable results with other SotA approaches and generalizes well to other frequencies where others fail.
In this section, we present a detailed comparison of our proposed methods, S4D-ViT-B and S5-ViT-B, with the SotA approaches in the domain of event-based vision for object detection, which can be seen in Table 1. The evaluation is conducted on two distinct datasets: Gen1 and 1 Mpx, to assess the performance and robustness of these methods under varying conditions.
@@
Our S4D-ViT-B and S5-ViT-B models, which integrate SSMs with Attention (3.2), demonstrate competitive performance across both datasets. Specifically, on the 1 Mpx dataset, our models achieve mAP scores of 46.8 and 47.2, respectively, which are competitive with the leading scores in this benchmark. While our models do not outperform the top-performing ERGO-12 [48] and GET-T [33] in terms of mAP, they exhibit a notable balance between accuracy and model complexity, as indicated by their parameter counts of 16.5M and 17.5M, respectively.
Notably, our models perform consistently well across different frequencies as it can be seen in Table 2, highlighting their robustness and generalizability in comparison to RVT and GET-T, which tend to exhibit large performance drops. We achieve an average performance drop between training and testing frequencies of 3.31 mAP averaged on both Gen1 [8] and 1 Mpx [34] datasets, while RVT [13] and GET [33] have a drop of 21.25 and 24.53 mAP, respectively.
In summary, our proposed S4D-ViT-B and S5-ViT-B models establish themselves as effective and efficient contenders in the field of event-based vision for object detection. Their competitive performance metrics, coupled with their generalization capabilities across various inference frequencies, make them valuable contributions to event-based vision community.
Model Dataset Frequency Evaluation (Hz) Perf. Drop 20 Hz 40 Hz 80 Hz 100 Hz 200 Hz RVT Gen1 47.16 35.13 21.98 18.61 8.35 26.14 1Mpx 47.40 42.51 33.20 30.29 18.15 16.36 S5 Gen1 47.40 46.44 45.08 42.49 39.84 3.94 1Mpx 47.20 46.49 46.11 45.80 39.70 2.68 GET Gen1 47.90 34.15 19.97 15.13 5.35 29.25 1Mpx 48.40 40.51 30.30 28.11 15.44 19.81
Table 2: Evaluation of RVT [13], S5 [41], and GET [33] across different frequencies on test datasets

### v3

@@ Abstract
Today, state-of-the-art deep neural networks that process event-camera data first convert a temporal window of events into dense, grid-like input representations. As such, they exhibit poor generalizability when deployed at higher inference frequencies (i.e., smaller temporal windows) than the ones they were trained on. We address this challenge by introducing state-space models (SSMs) with learnable timescale parameters to event-based vision. This design adapts to varying frequencies without the need to retrain the network at different frequencies. Additionally, we investigate two strategies to counteract aliasing effects when deploying the model at higher frequencies. We comprehensively evaluate our approach against existing methods based on RNN and Transformer architectures across various benchmarks, including Gen1 and 1 Mpx event camera datasets. Our results demonstrate that SSM-based models train 33% faster and also exhibit minimal performance degradation when tested at higher frequencies than the training input. Traditional RNN and Transformer models exhibit performance drops of more than 20 mAP, with SSMs having a drop of 3.76 mAP, highlighting the effectiveness of SSMs in event-based vision tasks.
Figure 1: Top-Left Previous works [13, 34] use RNN architectures with convolutional or attention mechanisms to train models that have superior performance on downstream tasks. However, the use of RNNs leads to slower training, and the learned weights only generalize well to data deployed at the same frequency as that used at training time. Bottom-Left We solve this problem by utilizing SSMs for temporal aggregation, which enables faster training by either utilizing the S4 model [19] or S5 [41] parallel scans. By their nature, these models allow deployment at different frequencies than those used at training time since they have a learnable timescale parameter. Right Our SSM-based models achieve an average performance drop between training and testing frequencies of 3.76 mAP averaged on both Gen1 [8] and 1 Mpx [34] datasets, while RVT [13] and GET [33] have a drop of 21.25 and 24.53 mAP, respectively.
@@ 1 Introduction
State-space models [16] function as CNN during training and are converted to an efficient RNN at test time. To achieve this, S4 [19] employs a technique where the SSM, which is not practical for training on modern hardware due to its sequential nature, is transformed into a CNN by unrolling. S5 [41] uses parallel scans during training, and is employed as RNN during inference. State-space models [16] can be deployed at different frequencies at inference time because they are Linear Time-Invariant (LTI) continuous-time systems that can be transformed into discrete-time system with an arbitrary step size. This feature permits inference at arbitrary frequencies by globally adjusting the timescale parameter based on the ratio between the new and old frequency sampling rates. Consequently, we tackle the longstanding issue in event-based vision, which requires multiple training cycles with different frequencies to adapt the neural network for various frequencies during inference.
For the task of object detection, we find that incorporating state-space layers accelerates the training by up to 33% relative to existing recurrent vision transformer approaches [13, 29]. This is achieved while maintaining performance comparable to existing methods. Notably, our approach demonstrates superior generalization to higher temporal frequencies with a drop of only 3.76 mAP, while previous methods experience a drop of 21.25 mAP or more. Also, we achieve comparable performance to RVT although we use a linear state-space model rather than a non-linear LSTM model. This also shows that the complexity of the LSTM might not be needed. For this to work, we introduce two strategies (frequency-selective masking and $H_{2}$ norm) to counteract the aliasing effect encountered with increased temporal frequencies in event cameras. The first one is a low-pass bandlimiting modification to SSM that encourages the learned convolutional kernels to be smooth. The second one mitigates the aliasing problem by attenuating the frequency response after a chosen frequency. We argue that state-space modeling offers a significant new direction for research in event-based vision, offering promising solutions to the challenges inherent in processing event-based data effectively and efficiently.
@@
  • Our experimental results outperform existing methods at higher frequencies by 20 mAP on average and show a 33% increase in the training speed.
@@ 2.1 Object detection with Event Cameras
The second research branch investigates the use of Graph Neural Networks (GNNs) or Spiking Neural Networks (SNNs). GNNs dynamically construct a spatio-temporal graph where new nodes and edges are instantiated by selectively sub-sampling events and identifying pre-existing nodes that are proximate in the space-time continuum [12, 43, 23]. A key challenge lies in architecting the network such that information can be disseminated effectively across extensive regions of the space-time volume. This becomes particularly important when dealing with large objects that exhibit slow relative motion to the camera. Moreover, while aggressive sub-sampling is often necessary to achieve low-latency inference, it introduces the risk of omitting important information from the event stream. On the other hand, SNNs [39, 7, 2] transmit information sparsely within the system. Unlike RNNs, SNNs emit spikes only when a threshold is met, making them hard to optimize due to the non-differentiable spike mechanism. Some solutions [32] bypass the threshold, but this sacrifices sparsity in deeper layers. Overall, SNNs remain a challenging area needing more foundational research for optimized performance.
@@ 2.2 Continuous-time Models
In our study, we empirically show that S4, S4D, and S5 models achieve results on par with state-of-the-art when combined with attention mechanisms on complex data. We also identify and address aliasing issues in these models, proposing two corrective strategies. Our work extends the range and robustness of continuous-time models for complex sequence modeling tasks such as object detection for event cameras.
@@ 3 Method
In this section, we first formalize the operating mechanism of event cameras and provide a notation used for describing state-space models in the preliminaries (Sec. 3.1). Secondly, we describe our approach of incorporating variants of state-space models as layers within our block (Sec. 3.2). This innovative design solves the problems associated with slow training and the variable frequency inference for event cameras.
@@ 3.1 Preliminaries
State-Space Models (SSMs). Linear State-Space Models (SSMs) form the crucial part of the backbone of our architecture, where we compare S4 [19], S4D [18], and S5 [41] layer variants. Given an input vector $\mathbf{u}(t)\in\mathbb{R}^{U}$, a latent state vector $\mathbf{x}(t)\in\mathbb{R}^{P}$, and an output vector $\mathbf{y}(t)\in\mathbb{R}^{M}$, the governing equations of a continuous-time linear SSM can be mathematically represented as:
@@ 3.2 SSM-ViT block
Regarding the block structure, initially, the input undergoes convolution with a defined kernel of size $k\times k$ with a stride $s$. This operation effectively captures the essential spatial features of the input. Following the convolution operation, the structure introduces a ’Block-SA’ module. This module is crucial in implementing self-attention mechanisms, but it does so specifically within local windows. The localized nature of the attention in this block ensures a focus on immediate spatial relations, allowing for a detailed representation of close-proximity features.
@@
The final and crucial component of the block structure is the State-Space Model (SSM) block, which is designed to compute both the output and the state in parallel, either by using S4 [19], S4D [18] or S5 [41] model variant. This ensures temporal consistency and a seamless transition of information between consecutive time steps. Efficient computation is crucial since it allows for faster training than RNN, and the timescale parameter on temporal aggregation is important since we can rescale it during inference and deploy it at any frequency we want.
@@ 3.3.1 Output Masking
To modulate the frequency spectrum of the model, we define a hyperparameter $\alpha$, which is used for masking the computed effective frequency $f_{n}$ for each state:
@@
where $\Delta t$ denotes the discrete time-step, and $r$ is the rate at which we train the model, by default it is 1, the rate gets halved when deploying at twice the trained frequency, etc.
@@
where $C_{n}$ represents the coefficients in the state representation. The selection of $\alpha$ is critical, with $\alpha=1.0$ representing the Nyquist limit. However, practical considerations and model constraints typically necessitate a lower empirical threshold. Our findings suggest that setting $\alpha=0.5$ yields optimal outcomes for systems with diagonal state matrices, such as in S4D and S5 configurations.
@@ 4 Experiments
We perform ablation studies and systematic evaluations of our proposed models utilizing both the Gen1 [8] and 1 Mpx [34] event-based camera datasets. We assess the model’s proficiency in adapting to unseen inference frequencies during training time on both datasets. We employed two variants for training across both datasets: the base model ViT-SSM-B, alongside its scaled-down derivative ViT-SSM-S (small). Additionally, to study the robustness and generalization capabilities of our architecture, we subject it to empirical testing on the DSEC dataset [14], which was not part of the original training corpus, and the visual results are provided in the appendix.
@@ 4.1 Setup
Implementation details. Our models are trained using 32-bit precision arithmetic across 400k iterations, making use of the ADAM optimizer [25] and a OneCycle learning rate schedule [42] that decays linearly from its highest value. We adopt a mixed batching technique that applies standard Backpropagation Through Time (BPTT) to half of the batch samples and Truncated BPTT (TBPTT) to the other half. This technique was introduced in the RVT [13]. We found it to be equally effective also for SSMs. For the S4(D) method, details can be found in [19, 18]. As for S5 [41], we utilize an efficient implementation from scratch in PyTorch; the original public version of S5 is available in JAX. Data augmentation is carried out through random horizontal flips and zooming operations, both inward and outward. Event representations are formed based on 50 ms time windows that correspond to 20 Hz sampling frequency, divided into T = 10 discrete bins. We incorporate a YOLOX detection head [11] that integrates IOU, class, and regression losses, averaged across both batch and sequence lengths during each optimization phase. Bandlimiting is implemented during training and inference of the SSM model by masking the output matrix C as explained in 3.3.1. In the case of $H_{2}$ norm, we add it as a term to the loss function of the model. We conduct training on the GEN1 dataset using A100 GPU, employing a batch size of 8 and a sequence length of 21. We use a global learning rate of 2e-4, which is propagated to the SSM components. For the 1 Mpx dataset, we train with a batch size of 12 and a sequence length of 10, using a learning rate of 3.5e-4 across two A100 GPUs.
@@
The 1 Mpx dataset [34] similarly focuses on driving environments but offers recordings at a higher resolution of 720 $\times$ 1280 pixels, captured over multiple months during both day-time and night-time. It contains roughly 15 hours of event data and accumulates around 25 million bounding box labels, spread across three classes: cars, pedestrians, and two-wheelers. Following the original evaluation criteria, we eliminate bounding boxes with a side length under 20 pixels and a diagonal less than 60 pixels, while also reducing the input resolution to 640 $\times$ 360 pixels.
@@ 4.2 Benchmark comparisons
S5-ViT-B (ours) Transformer + SSM YOLOX 47.7 8.16 47.8 9.57 18.2
S5-ViT-S (ours) Transformer + SSM YOLOX 46.6 7.81 46.5 8.87 9.7
Table 1: Comparisons on test sets of Gen1 and 1 Mpx datasets (20 Hz). Best results in bold and second best underlined. A star suggests that this information was not directly available and estimated based on the publications. Runtime is measured in milliseconds for a batch size of 1. We used a T4 GPU for SSM-ViT to compare against indicated timings in prior work [34] on comparable GPUs (Titan Xp). Our model is the 3rd best on the Gen1 and 2nd best on the 1Mpx dataset in terms of downstream task performance, while having fastest inference and less parameters.
In this section, we present a detailed comparison of our proposed methods, S4D-ViT and S5-ViT, with the SotA approaches in the domain of event-based vision for object detection, which can be seen in Table 1. The evaluation is conducted on two distinct datasets: Gen1 and 1 Mpx datasets.
@@
Our S4D-ViT-B and S5-ViT-B models, which integrate SSMs with Attention (3.2), demonstrate competitive performance across both datasets. Specifically, on the 1 Mpx dataset, our models register mAP scores of 46.8 for S4D and 47.8 and 46.5 for the S5 Base and Small variants, respectively. These scores are competitive with the top performances in this benchmark, securing the second-highest task performance on the 1 Mpx dataset. While our models do not outperform the top-performing ERGO-12 [48] and GET-T [33] in terms of mAP on the Gen1 dataset, they are very close to GET-T [33] with a difference of only 0.2 mAP.
Notably, our models perform consistently well across different frequencies as can be seen in Table 2, highlighting their robustness and generalizability in comparison to RVT and GET-T, which tend to exhibit large performance drops. We achieve an average performance drop between training and testing frequencies of 3.76 mAP averaged on both Gen1 [8] and 1 Mpx [34] datasets, while RVT [13] and GET [33] have a drop of 21.25 and 24.53 mAP, respectively.
In summary, our proposed S4D-ViT-B and S5-ViT-B models establish themselves as effective and efficient contenders in the field of event-based vision for object detection. Their competitive performance metrics, coupled with their generalization capabilities across various inference frequencies, make them valuable contributions to the event-based vision community.
Model Dataset Frequency Evaluation (Hz) Perf. Drop 20 Hz 40 Hz 80 Hz 100 Hz 200 Hz RVT Gen1 47.16 35.13 21.98 18.61 8.35 26.14 1Mpx 47.40 42.51 33.20 30.29 18.15 16.36 S5 Gen1 47.71 46.44 45.08 42.49 39.84 4.25 1Mpx 47.80 46.49 46.11 45.80 39.70 3.27 GET Gen1 47.90 34.15 19.97 15.13 5.35 29.25 1Mpx 48.40 40.51 30.30 28.11 15.44 19.81
Table 2: Evaluation of RVT [13], S5 [41], and GET [33] across different frequencies on test datasets. The Performance Drop is calculated as the average difference between the original performance at 20 Hz and performances at higher frequencies.